# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

### v1

@@ E.3 Variant (c)
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive UACQR-P UACQR-S
ailerons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.00
airfoil 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01
@@
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.96 $\pm$ 0.00
computer 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.97 $\pm$ 0.00
concrete 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.96 $\pm$ 0.02
elevator 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.97 $\pm$ 0.00
energy_efficiency 0.95 $\pm$ 0.01 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.95 $\pm$ 0.01 0.97 $\pm$ 0.02 0.96 $\pm$ 0.02
insurance 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
@@
qsar 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01
@@
Table 17: Variant (c) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive UACQR-P UACQR-S
ailerons 0.208 $\pm$ 0.016 0.227 $\pm$ 0.018 0.203 $\pm$ 0.017 0.237 $\pm$ 0.024 0.229 $\pm$ 0.020 0.177 $\pm$ 0.014 0.207 $\pm$ 0.017
airfoil 0.201 $\pm$ 0.014 0.284 $\pm$ 0.016 0.223 $\pm$ 0.024 0.204 $\pm$ 0.019 0.214 $\pm$ 0.024 0.365 $\pm$ 0.022 0.391 $\pm$ 0.029
allstate 0.278 $\pm$ 0.052 0.317 $\pm$ 0.065 0.276 $\pm$ 0.047 0.299 $\pm$ 0.065 0.315 $\pm$ 0.059 0.285 $\pm$ 0.051 0.305 $\pm$ 0.055
ca_housing 0.314 $\pm$ 0.010 0.426 $\pm$ 0.005 0.346 $\pm$ 0.013 0.332 $\pm$ 0.011 0.400 $\pm$ 0.016 0.366 $\pm$ 0.007 0.405 $\pm$ 0.007
computer 0.080 $\pm$ 0.004 0.103 $\pm$ 0.001 0.089 $\pm$ 0.004 0.084 $\pm$ 0.004 0.091 $\pm$ 0.006 0.084 $\pm$ 0.003 0.101 $\pm$ 0.001
concrete 0.278 $\pm$ 0.029 0.319 $\pm$ 0.023 0.283 $\pm$ 0.027 0.273 $\pm$ 0.031 0.276 $\pm$ 0.028 0.373 $\pm$ 0.025 0.403 $\pm$ 0.022
elevator 0.127 $\pm$ 0.008 0.192 $\pm$ 0.010 0.137 $\pm$ 0.008 0.152 $\pm$ 0.008 0.144 $\pm$ 0.010 0.158 $\pm$ 0.011 0.200 $\pm$ 0.010
energy_efficiency 0.043 $\pm$ 0.007 0.137 $\pm$ 0.007 0.051 $\pm$ 0.005 0.045 $\pm$ 0.007 0.049 $\pm$ 0.005 inf $\pm$ nan 0.162 $\pm$ 0.014
insurance 0.397 $\pm$ 0.092 0.420 $\pm$ 0.051 0.460 $\pm$ 0.051 0.395 $\pm$ 0.096 0.423 $\pm$ 0.083 0.305 $\pm$ 0.046 0.311 $\pm$ 0.033
kin8nm 0.324 $\pm$ 0.011 0.414 $\pm$ 0.018 0.331 $\pm$ 0.013 0.329 $\pm$ 0.011 0.344 $\pm$ 0.012 0.446 $\pm$ 0.018 0.460 $\pm$ 0.019
miami_housing 0.088 $\pm$ 0.001 0.112 $\pm$ 0.002 0.113 $\pm$ 0.008 0.093 $\pm$ 0.002 0.123 $\pm$ 0.008 0.104 $\pm$ 0.003 0.113 $\pm$ 0.002
naval_propulsion 1.6e-03 $\pm$ 3.0e-05 4.1e-03 $\pm$ 8.4e-05 1.6e-03 $\pm$ 4.1e-05 1.7e-03 $\pm$ 5.4e-05 1.8e-03 $\pm$ 3.9e-05 1.7e-03 $\pm$ 8.2e-05 3.6e-03 $\pm$ 8.5e-05
parkinsons 0.250 $\pm$ 0.007 0.373 $\pm$ 0.012 0.281 $\pm$ 0.011 0.260 $\pm$ 0.008 0.285 $\pm$ 0.010 0.269 $\pm$ 0.020 0.313 $\pm$ 0.018
powerplant 0.157 $\pm$ 0.007 0.176 $\pm$ 0.008 0.161 $\pm$ 0.008 0.169 $\pm$ 0.006 0.164 $\pm$ 0.007 0.193 $\pm$ 0.007 0.202 $\pm$ 0.009
qsar 0.345 $\pm$ 0.116 0.464 $\pm$ 0.153 0.354 $\pm$ 0.117 0.395 $\pm$ 0.135 0.385 $\pm$ 0.131 0.410 $\pm$ 0.135 0.452 $\pm$ 0.152
sulfur 0.111 $\pm$ 0.009 0.124 $\pm$ 0.010 0.109 $\pm$ 0.013 0.124 $\pm$ 0.009 0.114 $\pm$ 0.013 0.115 $\pm$ 0.012 0.126 $\pm$ 0.012
superconductor 0.194 $\pm$ 0.020 0.250 $\pm$ 0.025 0.211 $\pm$ 0.030 0.231 $\pm$ 0.025 0.308 $\pm$ 0.035 0.220 $\pm$ 0.022 0.232 $\pm$ 0.026
@@
Table 18: Variant (c) QUANTILELOSS at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive UACQR-P UACQR-S
ailerons 9.6e-06 $\pm$ 3.0e-07 1.0e-05 $\pm$ 2.7e-07 9.9e-06 $\pm$ 3.6e-07 1.1e-05 $\pm$ 2.7e-07 1.2e-05 $\pm$ 3.7e-07 1.0e-05 $\pm$ 0.0e+00 1.0e-05 $\pm$ 0.0e+00
airfoil 0.107 $\pm$ 0.014 0.140 $\pm$ 0.008 0.129 $\pm$ 0.019 0.110 $\pm$ 0.013 0.130 $\pm$ 0.019 0.188 $\pm$ 0.013 0.200 $\pm$ 0.010
allstate 1.2e+02 $\pm$ 6.844 1.2e+02 $\pm$ 5.120 1.2e+02 $\pm$ 19.422 1.4e+02 $\pm$ 9.738 1.6e+02 $\pm$ 10.676 1.1e+02 $\pm$ 7.356 1.2e+02 $\pm$ 6.507
ca_housing 2.8e+03 $\pm$ 91.370 3.2e+03 $\pm$ 69.918 3.2e+03 $\pm$ 1.0e+02 2.9e+03 $\pm$ 87.120 3.6e+03 $\pm$ 1.2e+02 3.0e+03 $\pm$ 58.564 3.0e+03 $\pm$ 57.081
computer 0.137 $\pm$ 0.011 0.152 $\pm$ 0.013 0.160 $\pm$ 0.012 0.140 $\pm$ 0.012 0.164 $\pm$ 0.013 0.147 $\pm$ 0.006 0.151 $\pm$ 0.004
concrete 0.335 $\pm$ 0.031 0.374 $\pm$ 0.043 0.380 $\pm$ 0.058 0.336 $\pm$ 0.032 0.384 $\pm$ 0.056 0.413 $\pm$ 0.037 0.425 $\pm$ 0.022
elevator 1.3e-04 $\pm$ 3.2e-06 1.7e-04 $\pm$ 2.3e-06 1.5e-04 $\pm$ 3.3e-06 1.4e-04 $\pm$ 3.7e-06 1.6e-04 $\pm$ 3.1e-06 1.8e-04 $\pm$ 5.4e-06 1.9e-04 $\pm$ 4.9e-06
energy_efficiency 0.029 $\pm$ 0.007 0.068 $\pm$ 0.004 0.034 $\pm$ 0.010 0.030 $\pm$ 0.007 0.034 $\pm$ 0.010 inf $\pm$ nan 0.079 $\pm$ 0.006
insurance 4.2e+02 $\pm$ 80.635 3.9e+02 $\pm$ 33.700 4.0e+02 $\pm$ 52.178 4.4e+02 $\pm$ 65.348 4.1e+02 $\pm$ 41.712 3.4e+02 $\pm$ 46.073 3.3e+02 $\pm$ 34.606
kin8nm 6.7e-03 $\pm$ 1.8e-04 8.4e-03 $\pm$ 2.7e-04 6.8e-03 $\pm$ 1.9e-04 6.9e-03 $\pm$ 2.2e-04 7.4e-03 $\pm$ 2.8e-04 8.8e-03 $\pm$ 1.6e-04 9.0e-03 $\pm$ 1.6e-04
miami_housing 3.7e+03 $\pm$ 1.4e+02 4.1e+03 $\pm$ 1.1e+02 7.3e+03 $\pm$ 4.2e+02 3.7e+03 $\pm$ 1.2e+02 7.7e+03 $\pm$ 4.2e+02 4.3e+03 $\pm$ 1.3e+02 4.5e+03 $\pm$ 1.3e+02
naval_propulsion 4.0e-05 $\pm$ 1.1e-06 9.5e-05 $\pm$ 2.1e-06 4.9e-05 $\pm$ 1.6e-06 4.2e-05 $\pm$ 1.2e-06 5.7e-05 $\pm$ 1.2e-06 9.3e-05 $\pm$ 9.0e-06 8.1e-05 $\pm$ 3.0e-06
parkinsons 0.189 $\pm$ 0.006 0.249 $\pm$ 0.007 0.217 $\pm$ 0.008 0.191 $\pm$ 0.006 0.220 $\pm$ 0.008 0.180 $\pm$ 0.009 0.200 $\pm$ 0.010
powerplant 0.203 $\pm$ 0.012 0.213 $\pm$ 0.014 0.208 $\pm$ 0.012 0.211 $\pm$ 0.013 0.211 $\pm$ 0.012 0.226 $\pm$ 0.011 0.234 $\pm$ 0.010
qsar 0.048 $\pm$ 0.002 0.057 $\pm$ 0.003 0.051 $\pm$ 0.002 0.053 $\pm$ 0.002 0.055 $\pm$ 0.003 0.051 $\pm$ 0.003 0.054 $\pm$ 0.002
sulfur 1.8e-03 $\pm$ 6.2e-05 2.0e-03 $\pm$ 1.2e-04 2.7e-03 $\pm$ 1.6e-04 1.8e-03 $\pm$ 6.2e-05 2.9e-03 $\pm$ 1.7e-04 1.9e-03 $\pm$ 1.2e-04 2.0e-03 $\pm$ 1.5e-04
superconductor 0.488 $\pm$ 0.020 0.553 $\pm$ 0.018 0.569 $\pm$ 0.042 0.564 $\pm$ 0.023 0.792 $\pm$ 0.035 0.507 $\pm$ 0.017 0.527 $\pm$ 0.016

### v2

@@ F.3 Variant (c)
ailerons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
airfoil 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
@@
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
concrete 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
elevator 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
energy_efficiency 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.95 $\pm$ 0.01 0.97 $\pm$ 0.01 0.96 $\pm$ 0.01
insurance 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
@@
qsar 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
@@
Table 23: Variant (c) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
ailerons 0.208 $\pm$ 0.017 0.227 $\pm$ 0.017 0.202 $\pm$ 0.018 0.237 $\pm$ 0.024 0.229 $\pm$ 0.020 0.208 $\pm$ 0.016 0.208 $\pm$ 0.017
airfoil 0.199 $\pm$ 0.013 0.286 $\pm$ 0.016 0.223 $\pm$ 0.024 0.204 $\pm$ 0.019 0.214 $\pm$ 0.024 0.205 $\pm$ 0.018 0.202 $\pm$ 0.014
allstate 0.279 $\pm$ 0.052 0.317 $\pm$ 0.065 0.277 $\pm$ 0.047 0.299 $\pm$ 0.065 0.315 $\pm$ 0.059 0.274 $\pm$ 0.053 0.272 $\pm$ 0.051
ca_housing 0.315 $\pm$ 0.010 0.425 $\pm$ 0.005 0.345 $\pm$ 0.013 0.332 $\pm$ 0.011 0.400 $\pm$ 0.016 0.312 $\pm$ 0.008 0.311 $\pm$ 0.008
computer 0.080 $\pm$ 0.004 0.102 $\pm$ 0.001 0.089 $\pm$ 0.004 0.084 $\pm$ 0.004 0.091 $\pm$ 0.006 0.079 $\pm$ 0.004 0.078 $\pm$ 0.004
concrete 0.278 $\pm$ 0.029 0.315 $\pm$ 0.019 0.284 $\pm$ 0.027 0.273 $\pm$ 0.031 0.276 $\pm$ 0.028 0.279 $\pm$ 0.029 0.281 $\pm$ 0.029
elevator 0.127 $\pm$ 0.008 0.192 $\pm$ 0.010 0.137 $\pm$ 0.008 0.152 $\pm$ 0.008 0.144 $\pm$ 0.010 0.127 $\pm$ 0.007 0.127 $\pm$ 0.007
energy_efficiency 0.043 $\pm$ 0.007 0.138 $\pm$ 0.007 0.051 $\pm$ 0.005 0.045 $\pm$ 0.007 0.049 $\pm$ 0.005 0.047 $\pm$ 0.007 0.046 $\pm$ 0.007
insurance 0.397 $\pm$ 0.091 0.421 $\pm$ 0.049 0.459 $\pm$ 0.050 0.395 $\pm$ 0.096 0.423 $\pm$ 0.083 0.440 $\pm$ 0.092 0.475 $\pm$ 0.109
kin8nm 0.325 $\pm$ 0.012 0.415 $\pm$ 0.018 0.330 $\pm$ 0.014 0.329 $\pm$ 0.011 0.344 $\pm$ 0.012 0.328 $\pm$ 0.014 0.326 $\pm$ 0.013
miami_housing 0.088 $\pm$ 0.001 0.112 $\pm$ 0.002 0.113 $\pm$ 0.008 0.093 $\pm$ 0.002 0.123 $\pm$ 0.008 0.078 $\pm$ 0.002 0.078 $\pm$ 0.002
naval_propulsion 1.6e-03 $\pm$ 3.2e-05 4.1e-03 $\pm$ 8.9e-05 1.6e-03 $\pm$ 4.3e-05 1.7e-03 $\pm$ 5.4e-05 1.8e-03 $\pm$ 3.9e-05 1.6e-03 $\pm$ 1.9e-05 1.6e-03 $\pm$ 3.0e-05
parkinsons 0.250 $\pm$ 0.007 0.371 $\pm$ 0.011 0.281 $\pm$ 0.011 0.260 $\pm$ 0.008 0.285 $\pm$ 0.010 0.251 $\pm$ 0.008 0.248 $\pm$ 0.007
powerplant 0.157 $\pm$ 0.007 0.175 $\pm$ 0.008 0.161 $\pm$ 0.008 0.169 $\pm$ 0.006 0.164 $\pm$ 0.007 0.158 $\pm$ 0.007 0.158 $\pm$ 0.007
qsar 0.344 $\pm$ 0.117 0.464 $\pm$ 0.153 0.353 $\pm$ 0.117 0.399 $\pm$ 0.138 0.386 $\pm$ 0.131 0.352 $\pm$ 0.120 0.346 $\pm$ 0.118
sulfur 0.111 $\pm$ 0.009 0.124 $\pm$ 0.010 0.109 $\pm$ 0.014 0.124 $\pm$ 0.009 0.114 $\pm$ 0.013 0.105 $\pm$ 0.009 0.104 $\pm$ 0.009
superconductor 0.195 $\pm$ 0.020 0.250 $\pm$ 0.025 0.210 $\pm$ 0.029 0.235 $\pm$ 0.027 0.308 $\pm$ 0.035 0.193 $\pm$ 0.020 0.193 $\pm$ 0.020
@@
Table 24: Variant (c) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
ailerons 9.6e-06 $\pm$ 2.9e-07 1.0e-05 $\pm$ 2.7e-07 9.9e-06 $\pm$ 3.6e-07 1.1e-05 $\pm$ 2.7e-07 1.2e-05 $\pm$ 3.7e-07 9.6e-06 $\pm$ 2.9e-07 9.6e-06 $\pm$ 2.9e-07
airfoil 0.107 $\pm$ 0.014 0.140 $\pm$ 0.007 0.129 $\pm$ 0.019 0.110 $\pm$ 0.013 0.130 $\pm$ 0.019 0.108 $\pm$ 0.014 0.108 $\pm$ 0.014
allstate 1.2e+02 $\pm$ 6.874 1.2e+02 $\pm$ 4.960 1.2e+02 $\pm$ 19.401 1.4e+02 $\pm$ 9.737 1.6e+02 $\pm$ 10.679 1.2e+02 $\pm$ 7.379 1.2e+02 $\pm$ 7.410
ca_housing 2.8e+03 $\pm$ 92.560 3.2e+03 $\pm$ 71.387 3.2e+03 $\pm$ 1.0e+02 2.9e+03 $\pm$ 87.120 3.6e+03 $\pm$ 1.2e+02 2.8e+03 $\pm$ 88.458 2.8e+03 $\pm$ 87.820
computer 0.137 $\pm$ 0.011 0.152 $\pm$ 0.013 0.160 $\pm$ 0.012 0.140 $\pm$ 0.012 0.164 $\pm$ 0.013 0.137 $\pm$ 0.011 0.138 $\pm$ 0.011
concrete 0.337 $\pm$ 0.032 0.374 $\pm$ 0.044 0.381 $\pm$ 0.059 0.336 $\pm$ 0.032 0.384 $\pm$ 0.056 0.336 $\pm$ 0.032 0.337 $\pm$ 0.032
elevator 1.3e-04 $\pm$ 3.2e-06 1.7e-04 $\pm$ 2.2e-06 1.5e-04 $\pm$ 3.3e-06 1.4e-04 $\pm$ 3.7e-06 1.6e-04 $\pm$ 3.1e-06 1.3e-04 $\pm$ 2.9e-06 1.3e-04 $\pm$ 2.9e-06
energy_efficiency 0.029 $\pm$ 0.007 0.068 $\pm$ 0.005 0.034 $\pm$ 0.010 0.030 $\pm$ 0.007 0.034 $\pm$ 0.010 0.030 $\pm$ 0.007 0.030 $\pm$ 0.007
insurance 4.2e+02 $\pm$ 80.269 3.9e+02 $\pm$ 33.576 4.0e+02 $\pm$ 53.283 4.4e+02 $\pm$ 65.348 4.1e+02 $\pm$ 41.712 4.3e+02 $\pm$ 77.643 4.5e+02 $\pm$ 67.994
kin8nm 6.7e-03 $\pm$ 1.8e-04 8.4e-03 $\pm$ 2.6e-04 6.8e-03 $\pm$ 1.9e-04 6.9e-03 $\pm$ 2.2e-04 7.4e-03 $\pm$ 2.8e-04 6.7e-03 $\pm$ 1.6e-04 6.7e-03 $\pm$ 1.7e-04
miami_housing 3.7e+03 $\pm$ 1.4e+02 4.2e+03 $\pm$ 1.2e+02 7.3e+03 $\pm$ 4.2e+02 3.7e+03 $\pm$ 1.2e+02 7.7e+03 $\pm$ 4.2e+02 4.0e+03 $\pm$ 2.4e+02 4.0e+03 $\pm$ 2.3e+02
naval_propulsion 4.0e-05 $\pm$ 1.0e-06 9.5e-05 $\pm$ 2.1e-06 4.9e-05 $\pm$ 1.6e-06 4.2e-05 $\pm$ 1.2e-06 5.7e-05 $\pm$ 1.2e-06 4.1e-05 $\pm$ 1.2e-06 4.5e-05 $\pm$ 1.6e-06
parkinsons 0.189 $\pm$ 0.006 0.249 $\pm$ 0.007 0.216 $\pm$ 0.008 0.191 $\pm$ 0.006 0.220 $\pm$ 0.008 0.189 $\pm$ 0.006 0.190 $\pm$ 0.007
powerplant 0.203 $\pm$ 0.012 0.213 $\pm$ 0.014 0.208 $\pm$ 0.012 0.211 $\pm$ 0.013 0.211 $\pm$ 0.012 0.203 $\pm$ 0.012 0.203 $\pm$ 0.012
qsar 0.048 $\pm$ 0.002 0.057 $\pm$ 0.003 0.051 $\pm$ 0.002 0.054 $\pm$ 0.002 0.055 $\pm$ 0.003 0.049 $\pm$ 0.002 0.049 $\pm$ 0.002
sulfur 1.8e-03 $\pm$ 6.3e-05 1.9e-03 $\pm$ 1.2e-04 2.7e-03 $\pm$ 1.6e-04 1.8e-03 $\pm$ 6.2e-05 2.9e-03 $\pm$ 1.7e-04 1.8e-03 $\pm$ 6.8e-05 1.8e-03 $\pm$ 7.1e-05
superconductor 0.489 $\pm$ 0.020 0.553 $\pm$ 0.018 0.568 $\pm$ 0.042 0.568 $\pm$ 0.025 0.792 $\pm$ 0.035 0.490 $\pm$ 0.020 0.489 $\pm$ 0.021